OpenStudy (anonymous):

Find if W1 is a linear subspace, being W1={(x;y;z) e R^2/ Sqrt(x+y+z)=0 } Find if W1 is a linear subspace, being W1={(x;y;z) e R^2/ Sqrt(x+y+z)=0 } @MIT 18.06 Linear …

OpenStudy (anonymous):

I supposse you mean \[W1 = \{ (x,y,z) \in R^3 ; \sqrt{x+y+z}=0\}\] which is the same set that a plane in 3D given by \[W2 = \{(x,y,z); x+y+z=0\}\]

OpenStudy (anonymous):

so because the plane is a subspace W1 is a subspace